OpenStudy (anonymous):

if x equals 5, then find the matching value for y. -x-y=-8

OpenStudy (anonymous):

-(5)-y=-8 y=-5+8 y=3

OpenStudy (anonymous):

thnks ;]

OpenStudy (bahrom7893):

Hey mathdrills, don't just give away the answers, teach them.

OpenStudy (anonymous):

Okay, you need to replace the x in the equation with the value of x you are given. -(5)-y=-8 Then, if you need help with the rest, just ask :-)

OpenStudy (bahrom7893):

lol better MathDrills =)